Acetylsalicylic acid is the active ingredient in aspirin. It took 35.17 ml of 0.5065 M sodium hydroxide to react completely with 3.210 grams of acetylsalicylic acid. Acetylsalicylic acid has one acidic hydrogen. What is the molecular mass of acetylsalicylic acid?
Answer #1

moles NaOH = M NaOH * LNaOH = ( 0.5065)(0.03517) = 0.0178 moles NaOH

moles NaOH = moles aspirin since aspirin has one acidic hydrogen

moles aspirin = ( 3.210 g / 180.157 g/mol) = 0.01781 moles

MM = ( g aspirin / moles aspirin) = ( 3.210 g / 0.01781 mol ) = 180.2 g / mol